# A Stabilized Unfitted Space-time Finite Element Method for Parabolic Problems on Moving Domains

### v1

@@ 1 Introduction
Recently, growing attention has been directed toward space-time FEMs that are continuous in both space and time [11, 12, 13, 14, 15]. These methods discretize the entire space-time domain using unstructured meshes and apply standard finite element techniques. A key advantage of this approach is its suitability for parallel computation over the full space-time domain, as well as its potential for space-time adaptive refinement. For instance, Steinbach et al. [11] introduced an unstructured conforming space-time FEM for general parabolic equations. Moore et al. [12] proposed a space-time isogeometric analysis using time-upwind test functions, applicable to both fixed and moving domains, which was later extended to multipatch isogeometric analysis in [13]. Localized stabilization strategies were further developed in [14, 15]. Additionally, a fitted space-time FEM was presented in [16] for advection-diffusion problems with moving interfaces. For a comprehensive overview of space-time FEMs for parabolic problems, we refer to [17]. However, when dealing with complex moving domain problems, the space-time methods mentioned above require the generation of a fitted and regular space-time mesh. This task becomes particularly challenging in three-dimensional spatial settings, where it entails constructing four-dimensional simplex meshes [18].
@@
Several unfitted FEMs have been developed for solving time-dependent problems involving moving domains or moving interfaces. In [32, 33], a time-stepping scheme is introduced for time-dependent moving domain problems, which requires a discrete extension of the solution from the previous time step to the domain at the next time step. Meanwhile, a space-time FEM based on discontinuous Galerkin (DG) time-stepping has been proposed in [34, 35, 36, 37] to handle moving domains and interfaces. This approach has also been adapted to moving surface problems [38] and coupled bulk-surface problems [39]. More recently, the method has been extended to a continuous-in-time space-time formulation [40, 41]. However, all the aforementioned methods are sequential in time and require sufficiently small time steps to maintain accuracy when the domain undergoes high oscillations or large deformations over time.
In this work, we propose a new unfitted space-time FEM that employs an unstructured mesh over the space-time domain along with piecewise linear finite elements. The use of unfitted meshes significantly simplifies mesh generation and reduces computational cost compared to fitted space-time approaches. A key advantage of our method over unfitted discontinuous Galerkin (DG) time-stepping schemes lies in its flexibility in space-time discretization. This flexibility enables local mesh refinement in regions where the domain undergoes rapid motion or large deformations, thereby facilitating fully adaptive resolution in both space and time while reducing the overall degrees of freedom. Furthermore, the resulting discrete system can be solved in a fully parallel manner via domain decomposition and efficient space-time preconditioning techniques, as exemplified in references [12, 42]. To mitigate the small cut issue, ghost penalty stabilization is employed to ensure a well-conditioned stiffness matrix. For handling convection-dominated cases, a streamline upwind Petrov–Galerkin (SUPG) scheme, inspired by [12, 15], is applied to stabilize the time-advection term.
@@
The remainder of this paper is organized as follows. Section 2 introduces the model problem for a parabolic moving domain and presents its variational formulation. Section 3 details the proposed unfitted space-time FEM. The corresponding energy error analysis is provided in Section 4. In Section 5, we initially establish a space-time Poincaré–Friedrichs inequality under mild geometric assumptions, followed by deriving an estimate for the condition number. Numerical experiments to validate our theoretical findings are presented in Section 6. Finally, Section 7 concludes the paper.
@@ 2 Model problem
We begin by introducing notation for the moving domain. Let $\Omega(t)\subset\mathbb{R}^{d}$, with $d=1,2,3$, be an open bounded domain whose boundary $\partial\Omega(t)$ is Lipschitz (for $d=2,3$). Here, $t$ indicates the time dependence of the domain. For simplicity, we denote $\Omega(t)$ as $\Omega_{t}$ and $\partial\Omega(t)$ as $\partial\Omega_{t}$ when appropriate.
Let $T>0$ and $I=(0,T)$ be the time interval. The associated space-time domain is defined as $Q:=\Omega(t)\times I$. The boundary of $Q$ is composed of three parts: the lateral (moving) boundary $\Sigma_{s}:=\partial\Omega(t)\times I$, the initial boundary $\Sigma_{0}:=\Omega(0)\times\{0\}$, and the final boundary $\Sigma_{T}:=\Omega(T)\times\{T\}$. We assume that $\partial Q$ is Lipschitz continuous and piecewise $C^{2}$-smooth. Moreover, we denote the space gradient operator by $\nabla_{x}$ and the space-time gradient operator by $\nabla$, i.e., $\nabla_{x}=\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{2}},\cdots,\frac{\partial}{\partial x_{d}}\right)^{T},\nabla=\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{2}},\cdots,\frac{\partial}{\partial x_{d}},\frac{\partial}{\partial t}\right)^{T}$.
@@
$\begin{cases}u_{t}-\nabla_{x}\cdot(a(x,t)\nabla_{x}u)=f&\text{in}~Q,\\ u=g&\text{on}~\Sigma_{s},\\ u(x,0)=u_{0}&\text{on}~\Sigma_{0},\end{cases}$ (2.1)
where $f\in H^{-1,0}(Q)$, $g\in H^{\frac{1}{2}}(\Sigma_{s})$, $u_{0}\in H^{1}(\Omega_{0})$, with $u_{0}$ and $g$ satisfying the compatibility conditions. The time-dependent diffusion coefficient $a(x,t)\in C^{0}(\overline{Q})$ is uniformly bounded from below and above by some positive constants, namely,
@@
We work under the following mild geometric assumptions::
@@
    $B(\alpha_{i},\beta_{i})=\big\{(\xi,\tau)\in\mathbb{R}^{d+1}:-1<\xi_{k}<1,\ \alpha_{i}<\tau<\beta_{i},\ k=1,\dots,d\big\},$
@@
For a real number $r\geq 0$ and a measurable set $U\subset\mathbb{R}^{d+1}$, let $H^{r}(U)$ denote the Sobolev space equipped with the usual norm $\left\|\cdot\right\|_{r,U}$ and semi-norm $\left|\cdot\right|_{r,U}$. By convection, for $r=0$, we write $\left(\cdot,\cdot\right)_{U}$ and $\left\|\cdot\right\|_{U}$ for the $L^{2}(U)$ inner product and norm, omitting the superscript $r$.
@@
To establish the weak formulation, we introduce additional function spaces, whose definitions and properties can be found in [45]. For $u\in H_{s}^{1,0}(Q)$, the distributional time derivative $\partial_{t}u$ is defined as the linear functional
@@
$\mathcal{W}_{s}:=\left\{\phi\in C^{1}(\overline{Q}):\phi~\text{is zero in a neighborhood of}~\Sigma_{s}\right\}.$
Having $\mathcal{W}_{s}\subset H_{s}^{1,0}(Q)$ from (2.2), and noticing from [2, Lemma 10.3.1] that $\partial_{t}\phi\in H^{-1,0}(Q)$ for all $\phi\in\mathcal{W}_{s}$, we define a norm on $\mathcal{W}_{s}$ by
$\|u\|_{W}^{2}=\|u\|_{H^{1,0}(Q)}^{2}+\|\partial_{t}\phi\|_{H^{-1,0}(Q)}^{2},\quad\forall u\in\mathcal{W}_{s}.$
Let $W_{s}$ be the closure of $\mathcal{W}_{s}$ in $H_{s}^{1,0}(Q)$ under $\|\cdot\|_{W}$. Then, under Assumption A.2,
$W_{s}=\left\{u\in H_{s}^{1,0}(Q):\partial_{t}u\in H^{-1,0}(Q)\right\}.$
In addition, a continuous linear trace operator can be defined from $W_{s}$ to each time slice $\Sigma_{\tau}:=\Omega_{\tau}\times\left\{t=\tau\right\}$ for $\tau\geq 0$. We define the subspace of $W_{s}$ consisting of functions with vanishing initial trace:
$W_{s,0}:=\left\{u\in W_{s}:u|_{\Sigma_{0}}=0\right\}.$
It is clear that $W_{s,0}\subset W_{s}\subset H_{s}^{1,0}(Q)$. We also introduce the broader space
$W=\left\{u\in H^{1,0}(Q):\partial_{t}u\in H^{-1,0}(Q)\right\}.$
@@
$A(u,v):=\left\langle u_{t},v\right\rangle+\left(a\nabla_{x}u,\nabla_{x}v\right)_{Q}=\left\langle f,v\right\rangle,\quad\forall v\in H_{s}^{1,0}(Q).$ (2.3)
The well-posedness of the variational formulation (2.3) follows by homogenizing the boundary conditions [47] and applying [45, Lemma 3.1].
@@
Suppose Assumption A.2 holds. Then for every $f\in H^{-1,0}(Q)$, $g\in H^{\frac{1}{2}}(\Sigma_{s})$, $u_{0}\in H^{1}(\Omega_{0})$, there exists a unique solution $u\in W$ to problem (2.3).
@@ 3 Unfitted space-time FEM
In this section, we discretize the variational formulation (2.3) using an unfitted space-time FEM. We begin by introducing some necessary notation. Let $\widetilde{\Omega}\subset\mathbb{R}^{d}$ be a polygonal background domain that contains the physical domain $\Omega(t)$ for all $t\in I$, i.e., $\Omega(t)\subset\widetilde{\Omega}$, and define the background space-time domain as $\widetilde{Q}=\widetilde{\Omega}\times I$. Let $\mathcal{T}_{h}$ be a family of conforming, quasi-uniform and regular triangulations of $\widetilde{Q}$, which are not required to align with the moving boundary $\Sigma_{s}$. Each element $K\in\mathcal{T}_{h}$ is considered as closed. Define the local mesh size $h_{K}=\text{diam}(K)$ and the global mesh size $h=\max_{K\in\mathcal{T}_{h}}h_{K}$. We introduce the so-called active mesh
@@
For every $K\in\mathcal{T}_{h}^{cut}$, there exists an element $K^{\prime}\in\mathcal{T}_{h}^{int}$ and a path $\{K_{1},K_{2},\dots,K_{N}\}\subset\mathcal{T}_{h}^{act}$ such that $K_{1}=K,K_{N}=K^{\prime}$, and consecutive elements $K_{i}$ and $K_{i+1}$ share a common facet for $i=1,\dots,N-1$. The path length $N$ is uniformly bounded by a constant independent of both the mesh size $h$ and the position of $\partial Q$ relative to $\mathcal{T}_{h}$.
@@
Next, we introduce some notation related to mesh faces. Let $\mathcal{F}_{h}$ denote the set of all element faces in $\mathcal{T}_{h}$, and define the following subsets:
@@
For any subcollection $\mathcal{T}\subset\mathcal{T}_{h}$ and any measurable set $U\subset\mathbb{R}^{d+1}$, we define
@@
where the continuous bilinear form $a_{h}$ and the linear form $l_{h}$ are defined as
$a_{h}(u,v)$= $(u_{t},v)_{Q}+(a\nabla_{x}u,\nabla_{x}v)_{Q}-\left\langle a\nabla_{x}u\cdot n_{x},v\right\rangle_{\Sigma_{s}}-\left\langle a\nabla_{x}v\cdot n_{x},u\right\rangle_{\Sigma_{s}}$ (3.2)
$+\left\langle\gamma h^{-1}u,v\right\rangle_{\Sigma_{s}}+\left\langle u,v\right\rangle_{\Sigma_{0}}+\left(\delta h^{2}(u_{t}-\nabla_{x}\cdot(a\nabla_{x}u)),v_{t}\right)_{\mathcal{T}_{h}^{act}\cap Q},$
$l_{h}(v):=\left\langle f,v\right\rangle-\left\langle a\nabla_{x}v\cdot n_{x},g\right\rangle_{\Sigma_{s}}+\left\langle\gamma h^{-1}g,v\right\rangle_{\Sigma_{s}}+\left\langle u_{0},v\right\rangle_{\Sigma_{0}}+\left(\delta h^{2}f,v_{t}\right)_{\mathcal{T}_{h}^{act}\cap Q},$
@@ 4 A priori error analysis
This section presents an a priori error estimate for the proposed method. Throughout the remainder of the paper, we assume that Assumptions A.1A.3 hold, unless stated otherwise. To this end, we define the energy space as $V=\left\{v\in W:v|_{K}\in H^{2}(K),~\forall K\in\mathcal{T}_{h}\right\}$. To facilitate the error estimates, we introduce the following norms and semi-norms. For any $v\in V_{h}+V$, define
$\left\|\thinspace\left|v\right|\thinspace\right\|^{2}$ $=\|a^{\frac{1}{2}}\nabla_{x}v\|_{Q}^{2}+\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v\|_{\Sigma_{s}}^{2}+\|v\|_{\Sigma_{0}}^{2}+\|v\|_{\Sigma_{T}}^{2}+\|\delta^{\frac{1}{2}}hv_{t}\|_{Q}^{2},$

### v2

@@ 1 Introduction
Recently, growing attention has been directed toward space-time FEMs that are continuous in both space and time [11, 12, 13, 14, 15, 16]. These methods discretize the entire space-time domain using unstructured meshes and apply standard finite element techniques. A key advantage of this approach is its suitability for parallel computation over the full space-time domain, as well as its potential for space-time adaptive refinement. For instance, Steinbach et al. [11] introduced an unstructured conforming space-time FEM for general parabolic equations. Moore et al. [12] proposed a space-time isogeometric analysis using time-upwind test functions, applicable to both fixed and moving domains, which was later extended to multipatch isogeometric analysis in [13]. Localized stabilization strategies were further developed in [14, 15]. Additionally, a fitted space-time FEM was presented in [16] for advection-diffusion problems with moving interfaces. For a comprehensive overview of space-time FEMs for parabolic problems, we refer to [17]. However, when dealing with complex moving domain problems, the space-time methods mentioned above require the generation of a fitted and regular space-time mesh. This task becomes particularly challenging in three-dimensional spatial settings, where it entails constructing four-dimensional simplex meshes [18].
@@
Several unfitted FEMs have been developed for solving time-dependent problems involving moving domains or moving interfaces. In [32, 33], a time-stepping scheme is introduced for time-dependent moving domain problems, which requires a discrete extension of the solution from the previous time step to the domain at the next time step. This idea has also been applied in [34] to solve two-phase flow problems with moving interfaces. In [35], a CutFEM for the one-phase Stefan–Signorini problem with applications in laser manufacturing is proposed. In [36], an interface-tracking algorithm combined with high-order unfitted characteristic finite element methods is developed for moving interface problems governed by the Oseen equations. Furthermore, several immersed finite element methods incorporating time-stepping strategies have been proposed in [37, 38, 39] for solving parabolic moving interface problems. Meanwhile, a space-time FEM based on discontinuous Galerkin (DG) time-stepping has been proposed in [40, 41, 42, 43] to handle moving domains and interfaces. This approach has also been adapted to moving surface problems [44] and coupled bulk-surface problems [45]. More recently, the method has been extended to a continuous-in-time space-time formulation [46, 47]. However, all the aforementioned methods are sequential in time and require sufficiently small time steps to maintain accuracy when the domain undergoes high oscillations or large deformations over time.
In this work, we propose a new unfitted space-time FEM that employs an unstructured mesh over the space-time domain along with piecewise linear finite elements. The use of unfitted meshes significantly simplifies mesh generation and reduces computational cost compared to fitted space-time approaches. A key advantage of our method over unfitted discontinuous Galerkin (DG) time-stepping schemes lies in its flexibility in space-time discretization. This flexibility enables local mesh refinement in regions where the domain undergoes rapid motion or large deformations, thereby facilitating fully adaptive resolution in both space and time while reducing the overall degrees of freedom. Furthermore, the resulting discrete system can be solved in a fully parallel manner via domain decomposition and efficient space-time preconditioning techniques, as exemplified in references [12, 48]. To mitigate the small cut issue, ghost penalty stabilization is employed to ensure a well-conditioned stiffness matrix. A streamline upwind Petrov–Galerkin (SUPG) scheme, inspired by [12, 15], is applied to stabilize the time-advection term.
@@
The remainder of this paper is organized as follows. Section 2 introduces the model problem posed on a moving domain and presents its variational formulation. Section 3 details the proposed unfitted space-time FEM. The corresponding energy error analysis is provided in Section 4. In Section 5, we initially establish a space-time Poincaré–Friedrichs inequality under mild geometric assumptions, followed by deriving an estimate for the condition number. Numerical experiments to validate our theoretical findings are presented in Section 6. Finally, Section 7 concludes the paper.
@@ 2 Model problem
We begin by introducing notation for the moving domain. Let $\Omega(t)\subset\mathbb{R}^{d}$, with $d=1,2,3$, be an open bounded domain whose boundary $\partial\Omega(t)$ is Lipschitz (for $d=2,3$). Here, $t$ indicates the time dependence of the domain. For simplicity, we denote $\Omega(t)$ as $\Omega_{t}$ and $\partial\Omega(t)$ as $\partial\Omega_{t}$ when appropriate. Moreover, we assume that the motion of the domain can be determined in advance based on prior knowledge–for instance, through a level set function $\phi(x,t)$ such that $\Omega(t)=\{x\in\mathbb{R}^{d}:\phi(x,t)<0\}$.
Let $T>0$ and $I=(0,T)$ be the time interval. The associated space-time domain is defined as $Q:=\bigcup_{t\in I}\Omega(t)\times\{t\}$. The boundary of $Q$ is composed of three parts: the lateral (moving) boundary $\Sigma_{s}:=\bigcup_{t\in I}\partial\Omega(t)\times\{t\}$, the initial boundary $\Sigma_{0}:=\Omega(0)\times\{0\}$, and the final boundary $\Sigma_{T}:=\Omega(T)\times\{T\}$. We assume that $\partial Q$ is Lipschitz continuous and piecewise $C^{2}$-smooth. Moreover, we denote the space gradient operator by $\nabla_{x}$ and the space-time gradient operator by $\nabla$, i.e., $\nabla_{x}=\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{2}},\cdots,\frac{\partial}{\partial x_{d}}\right)^{T},\nabla=\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{2}},\cdots,\frac{\partial}{\partial x_{d}},\frac{\partial}{\partial t}\right)^{T}$.
@@
$\begin{cases}\partial_{t}u-\nabla_{x}\cdot(a(x,t)\nabla_{x}u)=f&\text{in}~Q,\\ u=g&\text{on}~\Sigma_{s},\\ u(x,0)=u_{0}&\text{on}~\Sigma_{0},\end{cases}$ (2.1)
where $f\in L^{2}(Q)$, $g\in H^{\frac{1}{2}}(\Sigma_{s})$, $u_{0}\in H^{1}(\Omega_{0})$, with $u_{0}$ and $g$ satisfying the compatibility conditions. The time-dependent diffusion coefficient $a(x,t)\in C^{0}(\overline{Q})$ is uniformly bounded from below and above by some positive constants, namely,
@@
We work under the following mild geometric assumptions:
@@
    $B(\alpha_{i},\beta_{i})=\big\{(\xi,\tau)\in\mathbb{R}^{d+1}:-1<\xi_{k}<1,\ k=1,\dots,d,\ \alpha_{i}<\tau<\beta_{i}\big\},$
@@
For a real number $r\geq 0$ and a measurable set $U\subset\mathbb{R}^{d+1}$, let $H^{r}(U)$ denote the standard Sobolev space equipped with the usual norm $\left\|\cdot\right\|_{r,U}$ and semi-norm $\left|\cdot\right|_{r,U}$. By convection, for $r=0$, we write $\left(\cdot,\cdot\right)_{U}$ and $\left\|\cdot\right\|_{U}$ for the $L^{2}(U)$ inner product and norm, omitting the subscript $r$.
@@
To establish the weak formulation, we introduce additional function spaces, whose definitions and properties can be found in [51]. For $u\in H^{1,0}(Q)$, the distributional time derivative $\partial_{t}u$ is defined as the linear functional
@@
$W=\left\{u\in H^{1,0}(Q):\partial_{t}u\in H^{-1,0}(Q)\right\},$
with the norm
$\|u\|_{W}^{2}=\|u\|_{H^{1,0}(Q)}^{2}+\|\partial_{t}u\|_{H^{-1,0}(Q)}^{2}.$
@@
$A(u,v):=\left\langle\partial_{t}u,v\right\rangle+\left(a\nabla_{x}u,\nabla_{x}v\right)_{Q}=\left(f,v\right)_{Q},\quad\forall v\in H_{s}^{1,0}(Q).$ (2.3)
The well-posedness of the variational formulation (2.3) follows by homogenizing the boundary conditions [53] and applying [51, Lemma 3.1].
@@
Suppose Assumption A.2 holds. Then for every $f\in L^{2}(Q)$, $g\in H^{\frac{1}{2}}(\Sigma_{s})$, $u_{0}\in H^{1}(\Omega_{0})$, there exists a unique solution $u\in W$ to problem (2.3).
@@ 3 Unfitted space-time FEM
In this section, we discretize the variational formulation (2.3) using an unfitted space-time FEM. We begin by introducing some necessary notations. Let $\widetilde{\Omega}\subset\mathbb{R}^{d}$ be a polygonal background domain that contains the physical domain $\Omega(t)$ for all $t\in I$, i.e., $\Omega(t)\subset\widetilde{\Omega}$, and define the background space-time domain as $\widetilde{Q}=\widetilde{\Omega}\times I$. Let $\mathcal{T}_{h}$ be a family of conforming triangulations of $\widetilde{Q}$, which are not required to align with the moving boundary $\Sigma_{s}$. Each element $K\in\mathcal{T}_{h}$ is considered as closed. Define the local mesh size $h_{K}=\text{diam}(K)$ and the global mesh size $h=\max_{K\in\mathcal{T}_{h}}h_{K}$. We assume that $\mathcal{T}_{h}$ is shape regular and quasi-uniform, in the sense that there exist two constants $\mu,\nu>0$ such that each element $K\in\mathcal{T}_{h}$ contains a ball of diameter $\rho_{K}$ with $h_{K}/\rho_{K}\leq\mu$ and
$h/h_{K}\leq\nu,\quad\forall K\in\mathcal{T}_{h}.$ (3.1)
We introduce the so-called active mesh
@@
For every $K\in\mathcal{T}_{h}^{cut}$, there exists an element $K^{\prime}\in\mathcal{T}_{h}^{int}$ and a path $\{K_{1},K_{2},\dots,K_{N}\}\subset\mathcal{T}_{h}^{act}$ such that $K_{1}=K,K_{N}=K^{\prime}$, and consecutive elements $K_{i}$ and $K_{i+1}$ share a common facet for $i=1,\dots,N-1$. The path length $N$ is uniformly bounded by a constant independent of both the mesh size $h$ and the position of $\Sigma_{s}$ relative to $\mathcal{T}_{h}$.
@@
Next, we introduce some notations related to mesh faces. Let $\mathcal{F}_{h}$ denote the set of all element faces in $\mathcal{T}_{h}$, and define the following subsets:
@@
For any subcollection $\mathcal{T}\subset\mathcal{T}_{h}$, any measurable set $U\subset\mathbb{R}^{d+1}$ and any norm $\left\|\cdot\right\|$, we define
@@
Analogous definitions apply when the subscripts $\mathcal{T}\cap U$ and $K\cap U$ are substituted by $\mathcal{T}$ and $K$, respectively.
@@
where the bilinear form $a_{h}$ and the linear form $l_{h}$ are defined as
$a_{h}(u,v)$= $(\partial_{t}u,v)_{Q}+(a\nabla_{x}u,\nabla_{x}v)_{Q}-\left\langle a\nabla_{x}u\cdot n_{x},v\right\rangle_{\Sigma_{s}}-\left\langle a\nabla_{x}v\cdot n_{x},u\right\rangle_{\Sigma_{s}}$ (3.3)
$+\left\langle\gamma h^{-1}u,v\right\rangle_{\Sigma_{s}}+\left\langle u,v\right\rangle_{\Sigma_{0}}+\left(\delta h^{2}(\partial_{t}u-\nabla_{x}\cdot(a\nabla_{x}u)),\partial_{t}v\right)_{\mathcal{T}_{h}^{act}\cap Q},$
$l_{h}(v):=\left(f,v\right)_{Q}-\left\langle a\nabla_{x}v\cdot n_{x},g\right\rangle_{\Sigma_{s}}+\left\langle\gamma h^{-1}g,v\right\rangle_{\Sigma_{s}}+\left\langle u_{0},v\right\rangle_{\Sigma_{0}}+\left(\delta h^{2}f,\partial_{t}v\right)_{\mathcal{T}_{h}^{act}\cap Q},$
@@
The SUPG term is scaled with $h^{2}$ in the space-time mesh since we have assumed the shape regularity and quasi-uniformity of $\mathcal{T}_{h}$. This mesh parameter can be localized to $h_{K}^{2}$ if $\mathcal{T}_{h}$ is shape regular but not quasi-uniform. Analogous parameter selection is also valid for the terms $\left\langle\gamma h^{-1}u,v\right\rangle_{\Sigma_{s}}$ and $g_{h}(u,v)$, where the mesh parameter $h$ can be replaced by $h_{K}$ and $h_{F}:=\text{diam}(F)$, respectively. Moreover, the penalty parameters $\gamma$ and $\gamma_{1}$ can also be localized.
Remark 3.4.
@@ 4 A priori error analysis
This section presents an a priori error estimate for the proposed method. Throughout the remainder of the paper, we assume that Assumptions A.1A.3 hold, unless stated otherwise. To this end, we define the energy space as $V=\left\{v\in H^{1}(Q):v|_{K\cap Q}\in H^{2,1}(K\cap Q),~\forall K\in\mathcal{T}_{h}\right\}$. To facilitate the error estimates, we introduce the following norms and semi-norms. For any $v\in V_{h}+V$, define
$\left\|\thinspace\left|v\right|\thinspace\right\|^{2}$ $=\|a^{\frac{1}{2}}\nabla_{x}v\|_{Q}^{2}+\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v\|_{\Sigma_{s}}^{2}+\|v\|_{\Sigma_{0}}^{2}+\|v\|_{\Sigma_{T}}^{2}+\|\delta^{\frac{1}{2}}h\partial_{t}v\|_{Q}^{2},$